Find three consecutive positive odd integers such that the square of the smallest exceeds twice the largest by 91.
Thank you so much!

Answer by scott8148(6628)

x , x+2 , and x+4 are the integers

x^2 - 91 = 2(x + 4) ___ x^2 - 91 = 2x + 8

x^2 - 2x - 99 = 0

factoring ___ (x - 11)(x + 9) = 0